course: Control Systems.
please show all your work. I will rate you.

Find the value of K for the system given below that will place the closed-loop poles in the s-plane as shown. jw-axis
Answer #1

R(4) 130 / ou stept R®-voto hot R(S) - Feedback Slsto s 4512 Step 2 R() 152 C(s) 0 Slut) 875+2 feed bave feed bark = s(5+1) $K cseji SIS+Dr+s+2)t kl 1+1 feedback = k (5+) slsesez) tk k (Se 3+52 +24tk Step 3 R(S) as -- - ne (++) 33,5²25th the stret Ccoefficient y st must be zero therefore, Loefficient gs) = 2k zo K = 21*